Excel Online Structured Activity: Dividends Brooks Sporting Inc. is prepared to report the following 2016 income statement (sc. The company reported net income of $1.6 million in 2015. Assume that the number of shares outstanding has remained constanEXCEL SPREADSHEET DATA BELOW:

 
Dividends
Dollars in Thousands:
Tax rate 40.00%
Operating cost % 71.00%
Common shares outstanding 340,000
Common stock price $39.00
Dividend payout ratio 60.00%
Sales $11,800
Operating costs 8,378
EBIT $3,422
Interest 297
EBT $3,125
Taxes 1,250
Net income $1,875
Calculation of current per share dividend: Formulas
DPS, current year #N/A
Current dividend yield calculation:
Current dividend yield #N/A
Calculation of last year's per share dividend:
Last year's net income $1,600
DPS last year #N/A
Calculation of dividend payout ratio based on last year's per share dividend:
Dividend payout ratio on current net income #N/A
Answer #1

a). 2016 Net income = 1,875,000

Shares O/S = 340,000

Total dividend paid out in 2016 = payout ratio*net income = 60%*1,875,000 = 1,125,000

Dividend per share = 1,125,000/340,000 = $3.31

b). Current dividend yield = 2016 dividend/current share price = 3.31/39 = 8.48%

c). Dividend per share in 2015 = (60%*Net income/shares O/S) = 60%*1,600,000/340,000 = $2.82 per share

d). Same dividend per share in 2016 as 2015 means that total dividend paid out = 60%*1,600,000 = 960,000

Dividend payout ratio in 2016 = dividends paid out/net income = 960,000/1,875,000 = 51.20%

e). Same per-share dividend requires less capital than dividend based on same payout ratio so if the company wants to conserve capital, it is best to give the same dividend per share as in 2015. Statement II is correct.